OpenStudy (anonymous):

Let v1 = (2, -6) and v2 = (-4, 7). Compute the unit vectors in the direction of lv1l and lv2l

OpenStudy (psymon):

So to get unit vectors, we need to divide the vector by it's magnitude. So we would be doing: \[\frac{ 2i-6j }{ magnitude }\]and \[\frac{ -4i+7j }{ magnitude }\] Do you know how to compute the magnitude of a vector?

OpenStudy (anonymous):

I got 2root10 and root5?

OpenStudy (anonymous):

root65 for the second one actually

OpenStudy (psymon):

First one is okay, 2nd one might have been a typo, but sqrt(65). So now just divide the vector by its magnitude: \[\frac{ 2i }{ 2\sqrt{10} }-\frac{ 6j }{ 2\sqrt{10} }\] and \[\frac{ -4i }{ \sqrt{65} }+\frac{ 7j }{ \sqrt{65} }\] The first one can be simplied, the 2nd one cant. But those would be your unit vectors :3

OpenStudy (anonymous):

ok thanks so much :))

OpenStudy (psymon):

Yep, np ^_^

OpenStudy (anonymous):

Refer to the Mathematica 9 attachment.